Two angles are supplementary. The measure of the second angle is 45 degrees less than two times the measure of the first angle. Find the measure of each angle.
Please show me the formula and the steps to answer it correctly
pewe@quixnet.net

Answer by Starz(4)

alright....you can use a equation to solve this question.
Since we know that the two angles are supplementary, they must equal 180 degrees.
Let x represent the first angle.
The second angle is 45 degrees less than two times the measure of the first angle....so.....the second angle is 2x-45
If you put the equation together you will get x+2x-45=180
Now, let's solve:
x+2x-45=180
x+2x=180+45
3x=225
x=225 divided by 3
x=75
Now that you have the degrees of the first angles, you can find out the degrees of the second angle.
second angle=2x-45
=2 multiplied by 75-45
=150-45
=105
Therefore, the first angle is 75 degrees and the second angle is 105 degrees.
(let's check it to be sure)
75+105=180....yay!!!!!